# Conflict-Based Search as a Protocol: A Multi-Agent Motion Planning Protocol for Heterogeneous Agents, Solvers, and Independent Tasks

### v1

@@ Abstract
Imagine the future construction site, hospital, office, or even sophisticated household with dozens of robots bought from different manufacturers. How can we enable these different systems to effectively move in a shared environment, given that each robot may have its own independent motion planning system? This work shows how we can get efficient collision-free movements between algorithmically heterogeneous agents by using Conflict-Based Search (Sharon et al. 2015) as a protocol. At its core, the CBS Protocol requires one specific single-agent motion planning API; finding a collision-free path that satisfies certain space-time constraints. Given such an API, CBS uses a central planner to find collision-free paths - independent of how the API is implemented. We show how this protocol enables multi-agent motion planning for a heterogeneous team of agents completing independent tasks with a variety of single-agent planners including: Heuristic Search (e.g., A*), Sampling Based Search (e.g., RRT), Optimization (e.g., Direct Collocation), Diffusion, and Reinforcement Learning.
@@ I Introduction
The rapid progress of robotics technology is lowering robots’ costs while enhancing capabilities, leading to a future in which multi-agent robotic systems will be ubiquitous. Such systems may operate across a wide range of domains, including construction, manufacturing, office environments, and even sophisticated household settings. Importantly, these multi-agent teams are unlikely to be homogeneous; rather, they will consist of robots produced by different manufacturers, each performing independent tasks. This raises a fundamental question: how can heterogeneous robots, each potentially equipped with its own proprietary motion planning system, coordinate their movements in a shared environment?
@@
To that end, our main contribution is showing how we can use Conflict-Based Search as a protocol to enable efficient collision-free movement. Conflict-Based Search (CBS) [2] is a foundational method from Multi-Agent Path Finding [3] that finds collision-free multi-agent paths. At its core, we view CBS as defining a protocol that requires one specific single-agent motion planning API; finding a path that avoids space-time locations. Given such an API, CBS uses a central planner to resolve collision between agents - independent of how the API is implemented. Thus, “CBS as a protocol” enables collision-free multi-agent motion planning using different single-agent planners. We demonstrate the effectiveness of this protocol by showing how we can effectively plan for a heterogeneous team with a variety of single-agent planners including: Heuristic Search (e.g., A*), Sampling Based Search (e.g., RRT), Optimization (e.g., Direct Collocation), Diffusion, and Reinforcement Learning. Additionally, the CBS Protocol enables each agent to complete different independent tasks (e.g., coverage, surveillance) instead of just start-goal tasks.
@@ I-A Contributions
(1) Readers focused on single-agent motion planning but not multi-agent motion planning; For these readers, the key contribution of this paper is showing how we can use their single-agent motion planners in multi-agent systems through the CBS Protocol via defining one single API. The hope is that the protocol more easily enables single-agent motion planning researchers to incorporate their planners in multi-agent systems.
(2) Readers working on real-world MAMP but are unfamiliar with CBS, or only familiar with CBS in its classical (gridworld) context; For these readers, the primary contribution is showing how the CBS Protocol extends beyond gridworld and can incorporate a wide range of diverse motion planners (e.g., optimization solvers) for realistic kinodynamic motion planning.
(3) Readers familiar with CBS / MAPF and its extensions past non-gridworld; For these readers, the main contribution is the protocol formulation using heterogeneous solvers at once as well as some intricacies of the single-agent planners in the CBS context (e.g., using an RL policy).
@@ II Related Work
We first describe what we call “algorithmically homogeneous” MAMP methods that plan for agents using the same method across all agents, even if the agents themselves are heterogeneous. All of these methods explicitly require access to the internals of the agents and typically extend a single-agent planning technique to multiple agents. We then describe “algorithmically heterogeneous” MAMPF methods that plan for agents which have different single-agent planners.
@@ II-A1 Search-Based Planning
Search-Based Planning techniques for multi-agent motion planning typically focus on a simplified problem set-up called Multi-Agent Path Finding (MAPF) [3]. In the “classical” MAPF set-up, the world is discretized into a graph where each vertex corresponds to a grid cell. Time is also discretized into timestep where each agent can move to an adjacent cell or wait at its current location at each timestep (these define edges between the corresponding vertices). Each agent occupies exactly one vertex/grid cell at a timestep. At their core, search-based MAPF solvers use variants of A* space-time solvers. Operator Decomposition [4] addresses the combinatorial increase in search size by incrementally generating successors by planning individual agents. M* [5] uses a similar approach and is able to plan for up to 100 2D agents.
Conflict-Based Search [2] is a foundational technique in MAPF. We describe CBS more in-depth in Section III-A. Conceptually, CBS is a two-level heuristic search method that iteratively detects and resolves conflicts. CBS utilizes a high-level search that searches over constraints and a low-level search the plans single-agent paths. CBS first plans each agent individually and detects conflicts (i.e., collisions). CBS then proceeds to iteratively resolve conflicts by adding constraints to specific agents and replanning until a collision-free solution is found. Although originally created for MAPF with single-agent A* solvers, CBS has been used with other single-agent solvers as we will see later in this section. Several methods use different variants of A* solvers to handle heterogeneous teams of car-like robots with kinematic constraints [6, 7] or manipulator arms [8].
@@ II-A2 Sampling-Based Planning
Multi-Agent RRT* [9] is an initial attempt that tries to plan for up to 10 2D agents by running RRT* on the combined state space. Multi-Robot Discrete RRT (MRdRRT) [10] is able to plan for up to a combined 60 degrees of freedom over various agents with different degrees of freedom. MRdRRT’s main insight is to plan over a discrete implicit graph and to modify the extend operation to move agents sequentially in a priority order.
[11] plans for a heterogeneous team by generating a probabilistic roadmap for each agent and using CBS. Safe Interval Continuous-space CBS [12] recently proposed a novel single-agent Safe Interval RRT* with CBS to plan for up to 100 2D agents in continuous space.
@@ II-A3 Optimization Methods
Optimal Reciprocal Collision Avoidance (ORCA) [13] is able to plan for 1000’s of 2D agents by having each agent independently solve a low-dimensional linear program based on the positions and velocities of nearby agents. [14] introduces using safety control barrier certificates via formulating and solving a quadratic programming problem. Recently introduced Space-Time Graphs of Convex Sets [15] utilizes agent priorities and convex optimization to find solutions for up to 10 2D agents.
[16] first uses discrete planning (A*) in CBS and then post-processes the solution with optimization to plan for a heterogeneous drone and ground robot team. Discontinuity-Bound Conflict-Based Search (db-CBS) [17] interleaves discrete search and optimization by using a discontinuity-bounded A* search (e.g., A* with motion primitives) in CBS to generate single-agent approximate solutions which are used as warm-starts for a joint optimization (across all agents) that finds kinodynamically feasible solutions. If the optimization does not return a feasible solution, then constraints are added and the CBS search continues. Conflict-Based Model Predictive Control (CB-MPC) [18] replaces the low-level A* search in CBS with limited horizon MPC.
@@ II-A4 Machine Learning Methods
Multi-Agent DDPG [19] and Multi-Agent PPO [20] extrapolate single-agent DDPG [21] and PPO [22] respectively for multi-agent systems. These approaches are more general than collision-free MAMP but can be used for MAMP (e.g., [23]).
There are several works that try to use reinforcement learning or supervised learning for MAPF. Specifically, these works train a local individual agent policy that learns to avoid collisions [24] and can handle up to 10000 agents in congestion [25]. Recently, MMD [26] trained a diffusion model as a single-agent planner for CBS and incorporates constraints through guidance terms.
@@
Taking a step back, we note that CBS appears in every single category with different types of single-agent solvers, with their emphasis on the single-agent solver. CBS as a Protocol has a different focus and can be seen as encompassing all these methods. Our perspective of CBS using an abstracted single-agent API means that we can recover [2, 6, 7, 8, 11, 12, 16, 18, 26] with different API implementations.
@@ II-B Algorithmically Heterogeneous MAMP Techniques
Unlike algorithmically homogeneous MAMP methods, algorithmically heterogeneous MAMP methods do not require each agent run the same solver. To our surprise, we could not find any published work that addresses this problem setting. In particular, existing MAMP methods for heterogeneous teams focus on robots with different capabilities but use algorithmically homogeneous solutions (e.g., [7, 11, 16]). On the other hand, existing multi-agent task planning/coordination methods focus on heterogeneous behaviors or task assignment and not on collision-free movement [27, 28].
@@
To our knowledge, current solutions for algorithmically heterogeneous multi-robot teams in the real world currently plan robots individually viewing other robots as dynamic obstacles. The most common solution is for agents to plan independent paths and rely on local sensing to replan when another agent obstructs their path. However, this reduces their ability to move fast, requires additional planning time during execution, and can lead to deadlock. An alternative scheme could be to use Prioritized Planning [29] which assigns agents priorities and plans agents sequentially starting with the highest priority agent. Each agent plans a path that avoids colliding with the paths of previously planned (higher priority) agents. However, in congestion this can result in no solution being found, which limits its practicality. Taking a step back, the problem with both of these approaches is that they solve the multi-agent problem as greedy single-agent calls viewing other agents as obstacles, resulting in them failing in complex instances. To that end, we introduce CBS as a Protocol as a principled AH-MAMP protocol that does not view agents as obstacles but instead intelligently searches over space-time constraints to find collision-free solutions.
@@ III CBS as a Protocol
We first describe CBS from classical MAPF and then generalize CBS as a protocol.
@@ III-A Conflict Based Search
CBS is a two-level heuristic search algorithm for solving MAPF optimally. We focus on the top-level which searches over “Constraint Tree” (CT) nodes and describe its core search components (e.g., nodes, goal condition, etc). Algorithm 1 describes the CBS Protocol which generalizes CBS and can be used to follow along.
@@
3:   CTQueue.insert(node)
4:   while CTQueue.notEmpty() and not TimeOut() do
5:    node $\leftarrow$ CTQueue.pop()
@@
High-Level Search: CBS employs a best-first search over the CT nodes by using a priority queue (which we will call “CTQueue”). There are multiple ways to sort the priority queue based on the MAPF objective, but one standard way is to sort by CT node’s cost (typically the sum of each agent’s path length). CBS pops out the cheapest node and checks if it is a goal node (line 5). If it is, CBS returns that node’s paths as the solution. If not, CBS generates successor nodes and inserts them into CTQueue and repeats (line 17).
@@ III-B CBS as a Protocol
As mentioned in Sec II-A, although CBS has predominantly used A* single-agent planners, the general idea of iteratively planning and resolving collisions by searching over a Constraint Tree can be applied to other single-agent planners [17, 12, 18, 26]. However, those publications with CBS focus on developing the single-agent planner.
We come from the AH-MAMP perspective where we do not focus on single-agent solvers but instead want a mechanism to coordinate an algorithmically diverse set of agents. Our insight is that CBS can be used as a protocol across multiple different solvers at once. In particular, CBS solely requires each agent to have the ability to plan a path that satisfies the CT node’s constraints. Thus, each agent could run their own individual solver with their own kinematic constraints. Since different agents could have different cost functions (other than path length), each solver should also return the solution cost, which can be included in the CT node.
An important caveat is that given a set of agent plans, CBS must detect collisions between agents. Thus, we require the single-agent planner to not just return a path but also the space-time volume occupied by the agents path. Note that compact representations for this are possible (e.g., if the robot’s footprint is static, giving the footprint along with the space-time path defines the occupied space-time volume).
CBS as a protocol retains the exact same CBS structure with the main change of generalizing the single-agent planner. In particular, instead of the planner being predefined, we view the plan call as an API call that can be written independently.
@@
In the context of CBS, this API will be repeatedly queried with different constraints. Assuming the API is solving the same overall task between calls (e.g., start-goal task), this means the API is repeatedly solving the same problem with minor variations (through the constraints). Recent work [30] showed how CBS can be sped up by having its single-agent solver reuse “experience” from previous queries. This idea is broadly applicable to different types of single-agent motion planning solvers. Thus single-agent solver APIs are recommended to leverage experience from past queries for faster query times in the CBS Protocol.
Conflict Detection and Constraints: The CBS Protocol detects inter-agent collisions in a CT node’s solution by checking for overlaps between all pairs of agents’ space-time volume. If a collision is detected between two agents $A,C$, a representative space-time point $(p_{0},t_{0})$ in the overlapped region is selected (in 2D this corresponds to a point $p_{0}=(x_{0},y_{0})$). To resolve this collision, we add a $(\text{agent\_id},p_{0},t_{0})$ constraint for each colliding agent that forbids the agent from overlapping spatially with $(p_{0},t_{0})$ (we note prior work [31] introduced this constraint for large agents in gridworld MAPF). We highlight that the constraint is defined in the workspace dimension (e.g., 2D or 3D) and is independent of the dimensionality of each agent / solver.
A subtle but important caveat is that in continuous time motion planning, a constraint at a specific instantaneous time could be satisfied by planning a near identical path with an almost instantaneous wait, which will result in another collision at the same location at a negligibly offset time. Likewise for continuous space motion planning, a constraint at a single point could be satisfied by planning a near identical path with an almost negligible deviation. Thus, for practical purposes, the CBS Protocol should use constrain space-time volumes (e.g., forcing the agent to avoid $p_{0}$ from $(t_{0},t_{0}+\delta)$ where $\delta$ is the constraint duration) that force meaningful changes in agents’ paths. We note that there are extensive improvements in the CBS literature for designing better constraints [8, 32, 33], but this is not related to our main contribution and should be incorporated in the future.
Independent Tasks: CBS for motion planning has been primarily used to plan collision-free start-goal paths. From the API perspective, however, the API simply needs to find a solution that satisfies the constraints. A few prior works have used this flexibility in CBS for planning paths that visit multiple goal locations [34]. Thus more generally, the API is not limited to planning start-goal paths and can be used for arbitrary single-agent tasks (e.g., coverage, exploration).
@@
Theoretical Properties of the CBS Protocol: The theoretical guarantees of using the CBS Protocol depend on the properties of the single-agent APIs and the constraints used. The CBS Protocol is “complete” (i.e., guaranteed to eventually find a valid solution in a solvable instance) given two conditions: 1) Each single-agent API will find a valid solution with bounded cost that satisfies constraints if a solution exists, and 2) The constraints are mutually disjunctive. The proof of completeness follows regular CBS [2].
However, as mentioned earlier, for practical performance constraints should be on space-time volumes, which are not mutually disjunctive. More broadly, in continuous time planning, designing meaningful mutually disjunctive constraints for CBS is an unsolved question [35]. Thus, the CBS Protocol is not theoretically complete in the continuous space-time context, however, it still offers a principled way to resolve inter-agent conflicts for efficient motion planning.
IV Experimental Section

### v2

@@ Abstract
Imagine the future construction site, hospital, or office with dozens of robots bought from different manufacturers. How can we enable these different robots to effectively move in a shared environment, given that each robot may have its own independent motion planning system? This work shows how we can get efficient collision-free movements between algorithmically heterogeneous agents by using Conflict-Based Search (Sharon et al. 2015) as a protocol. At its core, the CBS Protocol requires one specific single-agent motion planning API; finding a collision-free path that satisfies certain space-time constraints. Given such an API, CBS uses a central planner to find collision-free paths - independent of how the API is implemented. We demonstrate how this protocol enables multi-agent motion planning for a heterogeneous team of agents completing independent tasks with a variety of single-agent planners including: Heuristic Search (e.g., A*), Sampling Based Search (e.g., RRT), Optimization (e.g., Direct Collocation), Diffusion, and Reinforcement Learning.
@@ I Introduction
The rapid progress of robotics technology is lowering robots’ costs while enhancing capabilities, leading to a future in which multi-agent robotic systems will be ubiquitous. Such systems may operate across a wide range of domains, including construction, manufacturing, office environments, and even sophisticated household settings. Importantly, these multi-agent teams are unlikely to be homogeneous; rather, they will consist of robots produced by different manufacturers, each performing independent tasks. This raises a fundamental question: how can heterogeneous robots, each potentially equipped with its own motion planning system, coordinate their movements in a shared environment?
@@
To that end, our main contribution is demonstrating how we can use Conflict-Based Search as a protocol to enable efficient collision-free movement. Conflict-Based Search (CBS) [sharon2015cbs] is a foundational method from Multi-Agent Path Finding [stern2019mapfbenchmark] that finds collision-free multi-agent paths. At its core, we view CBS as defining a protocol that requires one specific single-agent motion planning API; finding a path that avoids space-time locations. Given such an API, CBS uses a central planner to resolve collision between agents - independent of how the API is implemented. Thus, “CBS as a protocol” enables collision-free multi-agent motion planning using different single-agent planners. We demonstrate the effectiveness of this protocol by showing how we can effectively plan for a heterogeneous team with a variety of single-agent planners including: Heuristic Search, Sampling Based Search, Optimization, Diffusion, and Reinforcement Learning. Additionally, the CBS Protocol enables each agent to complete different independent tasks (e.g., coverage, surveillance) instead of just start-goal tasks.
@@ I-A Contributions
(1) Readers focused on single-agent motion planning but not MAMP; For these readers, the key contribution of this paper is showing how we can use their single-agent motion planners in multi-agent systems through the CBS Protocol via defining one single API. The hope is that the protocol more easily enables single-agent motion planning researchers to incorporate their planners into multi-agent systems.
(2) Readers working on real-world MAMP but are unfamiliar with CBS, or only familiar with CBS in its classical (gridworld) context; For these readers, the primary contribution is showing how the CBS Protocol extends beyond gridworld and can incorporate a range of diverse motion planners (e.g., optimization) for kinodynamic motion planning.
(3) Readers familiar with CBS and its extensions past non-gridworld; For these readers, the main contribution is the protocol formulation using heterogeneous solvers at once as well as some intricacies of the single-agent planners in the CBS context (e.g., using an RL policy). We note that the CBS Protocol high-level search does not differ from CBS.
@@ II Related Work
We first describe what we call “algorithmically homogeneous” MAMP methods that plan for agents using the same method across all agents, even if the agents themselves are heterogeneous. These methods explicitly require access to the internals of the agents and typically extend a single-agent planning technique to multiple agents. We then describe “algorithmically heterogeneous” MAMPF methods that plan for agents which have different single-agent planners.
@@ II-A1 Search-Based Planning
Search-Based Planning techniques for MAMP typically focus on a simplified problem set-up called Multi-Agent Path Finding (MAPF) [stern2019mapfbenchmark]. In the “classical” MAPF set-up, the world is discretized into a graph where each vertex corresponds to a grid cell. Time is also discretized into timestep where each agent can move to an adjacent cell or wait at its current location at each timestep (these define edges between the corresponding vertices). Each agent occupies exactly one vertex/grid cell at a timestep. Search-based MAPF solvers use variants of A* space-time solvers. Operator Decomposition [standley2010operater_decomposition] addresses the combinatorial increase in search size by incrementally generating successors by planning individual agents. M* [mstart_2011] uses a similar approach to plan for up to 100 2D agents.
Conflict-Based Search [sharon2015cbs] is a foundational technique in MAPF. We describe CBS more in-depth in Section III-A. Conceptually, CBS is a two-level heuristic search method that iteratively detects and resolves conflicts. CBS utilizes a high-level search that searches over constraints and a low-level search the plans single-agent paths. CBS first plans each agent individually and detects conflicts (i.e., collisions). CBS then proceeds to iteratively resolve conflicts by adding constraints to specific agents and replanning until a collision-free solution is found. Although originally created for MAPF with single-agent A* solvers, CBS has been used with other single-agent solvers as we will see later in this section. Several methods use different variants of A* solvers to handle heterogeneous teams of car-like robots with kinematic constraints [wen2022carlike_cbs, bai2025cbs_heterogeneous_robots] or manipulator arms [shaoul2024generalized_ecbs].
@@ II-A2 Sampling-Based Planning
Multi-Agent RRT* [vcap2013multi_agent_rrt] is an initial attempt that tries to plan for up to 10 2D agents by running RRT* on the combined state space. Multi-Robot Discrete RRT (MRdRRT) [solovey2016discrete_rrt] is able to plan for up to a combined 60 degrees of freedom over various agents with different degrees of freedom. MRdRRT’s main insight is to plan over a discrete implicit graph and to modify the extend operation to move agents sequentially in a priority order.
[solis2021representation_optimal_cbs] plans for a heterogeneous team by generating a probabilistic roadmap for each agent and using CBS. Safe Interval Continuous-space CBS [sim2024safe_interval_rrt] recently proposed a novel single-agent Safe Interval RRT* with CBS to plan for up to 100 2D agents in continuous space.
@@ II-A3 Optimization Methods
Optimal Reciprocal Collision Avoidance (ORCA) [alonso2013orca] is able to plan for 1000’s of 2D agents by having each agent independently solve a low-dimensional linear program based on the positions and velocities of nearby agents. [wang2017safety_barrier_certificates] introduces using safety control barrier certificates via formulating and solving a quadratic programming problem. Recently introduced Space-Time Graphs of Convex Sets [tang2025multiagent_space_time_gcs] utilizes agent priorities and convex optimization to find solutions for up to 10 2D agents.
[debord2018multi_drone_ground_cbs] first uses discrete planning (A*) in CBS and then post-processes the solution with optimization to plan for a heterogeneous drone and ground robot team. Discontinuity-Bound Conflict-Based Search (db-CBS) [moldagalieva2024dbcbs] interleaves discrete search and optimization by using a discontinuity-bounded A* search (e.g., A* with motion primitives) in CBS to generate single-agent approximate solutions which are used as warm-starts for a joint optimization across all agents. If the optimization does not return a feasible solution, then constraints are added and the CBS search continues. Conflict-Based Model Predictive Control (CB-MPC) [tajbakhsh2024cbmpc] replaces the low-level A* search in CBS with limited horizon MPC.
@@ II-A4 Machine Learning Methods
Multi-Agent DDPG [lowe2017maddpg] and Multi-Agent PPO [yu2022mappo] extrapolate single-agent DDPG [lillicrap2015ddpg] and PPO [schulman2017ppo] respectively for multi-agent systems. These approaches are more general than collision-free MAMP but can be used for MAMP (e.g., [peng2023mappohr]).
Several methods train a single agent policy for MAPF using reinforcement learning or supervise learning that learns to avoid collisions [sartoretti2019primal] and can handle up to 10000 agents in congestion [jiang2024sillm]. Other recent work MMD [shaoul2025mmd] trained a diffusion model as a single-agent planner for CBS and incorporates constraints through guidance terms.
@@
Taking a step back, we note that CBS appears in each category with different types of single-agent solvers, with their emphasis on the single-agent solver. CBS as a Protocol has a different focus and can be seen as encompassing all these methods. Our perspective of CBS using an abstracted single-agent API means that we can recover [sharon2015cbs, wen2022carlike_cbs, bai2025cbs_heterogeneous_robots, shaoul2024generalized_ecbs, solis2021representation_optimal_cbs, sim2024safe_interval_rrt, debord2018multi_drone_ground_cbs, tajbakhsh2024cbmpc, shaoul2025mmd] with different API implementations.
@@ II-B Algorithmically Heterogeneous MAMP Techniques
Unlike algorithmically homogeneous MAMP methods, algorithmically heterogeneous MAMP methods do not require each agent to run the same solver. To our surprise, we could not find any published work that addresses this problem setting. Existing MAMP methods for heterogeneous teams focus on robots with different capabilities but use algorithmically homogeneous solutions (e.g., [bai2025cbs_heterogeneous_robots, solis2021representation_optimal_cbs, debord2018multi_drone_ground_cbs]). On the other hand, existing multi-agent task planning/coordination methods focus on heterogeneous behaviors or task assignment and not on collision-free movement [rizk2019heterogeneous_multi_robot_survey, bettini2023heterogeneous_behavior_marl].
@@
To our knowledge, current solutions for algorithmically heterogeneous multi-robot teams in the real world currently plan robots individually viewing other robots as dynamic obstacles. The most common solution is for agents to plan independent paths and rely on local sensing to replan when another agent obstructs their path. However, this reduces their ability to move fast, requires additional planning time during execution, and can lead to deadlock. An alternative scheme could be to use Prioritized Planning [erdmann1987multiple] which assigns agents priorities and plans agents sequentially starting with the highest priority agent. Each agent plans a path that avoids colliding with the paths of previously planned (higher priority) agents. However, in congestion this can result in no solution being found, which limits its practicality. Taking a step back, the problem with both of these approaches is that they solve the multi-agent problem as greedy single-agent calls viewing other agents as obstacles, resulting in them failing in complex instances. To that end, we introduce CBS as a Protocol as a principled AH-MAMP protocol that does not view agents as obstacles but instead intelligently searches over space-time constraints to find collision-free solutions.
@@ III CBS as a Protocol
We first describe CBS from classical MAPF and then describe CBS as a protocol.
@@ III-A Conflict Based Search
CBS is a two-level heuristic search algorithm for solving MAPF optimally. We focus on the top-level which searches over “Constraint Tree” (CT) nodes and describe its core search components (e.g., nodes, goal condition, etc). Algorithm 1 describes the CBS Protocol (identical to CBS except for the plan API) and can be used to follow along.
@@
3:   OPEN.insert(node)
4:   while OPEN.notEmpty() and not TimeOut() do
5:    node $\leftarrow$ OPEN.pop()
@@
High-Level Search: CBS employs a best-first search over the CT nodes by using a priority queue (which is called “OPEN”). There are multiple ways to sort the priority queue based on the MAPF objective, but one standard way is to sort by CT node’s cost (typically the sum of each agent’s path length). CBS pops out the cheapest node and checks if it is a goal node (line 5). If it is, CBS returns that node’s paths as the solution. If not, CBS generates successor nodes and inserts them into OPEN and repeats (line 17).
@@ III-B CBS as a Protocol
As mentioned in Sec II-A, although CBS has predominantly used A* single-agent planners, the general idea of iteratively planning and resolving collisions by searching over a Constraint Tree can be applied to other single-agent planners [moldagalieva2024dbcbs, sim2024safe_interval_rrt, tajbakhsh2024cbmpc, shaoul2025mmd]. However, those publications with CBS focus on developing the single-agent planner.
We come from the AH-MAMP perspective where we do not focus on single-agent solvers but instead want a mechanism to coordinate an algorithmically diverse set of agents. Our insight is that CBS can be used as a protocol across multiple different solvers at once. CBS solely requires each agent to have the ability to plan a path that satisfies the CT node’s constraints. Thus, each agent could run their own individual solver with their own kinematic constraints. Since different agents could have different cost functions (other than path length), each solver should also return the solution cost, which can be included in the CT node.
An important caveat is that given a set of agent plans, CBS must detect collisions between agents. Thus, we require the single-agent planner to not just return a path but also the space-time volume occupied by it. Note that compact representations for this are possible (e.g., if the robot’s footprint is static, giving the footprint along with the space-time path defines the occupied space-time volume).
CBS as a protocol is exactly the same as CBS except for the change of perspective of the single-agent planner. In particular, instead of the planner being predefined, we view the plan call as an API call that can be written independently.
@@
In the context of CBS, this API will be repeatedly queried with different constraints. If the API is solving the same task between calls (e.g., start-goal task), this means the API is repeatedly solving the same problem with minor variations (through the constraints). [shaoul2024experience-cbs] showed how CBS can be sped up by having its single-agent solver reuse “experience” from previous queries. This idea is broadly applicable to different types of single-agent motion planning solvers. Thus single-agent solver APIs are recommended to leverage experience from past queries for faster query times in the CBS Protocol.
Conflict Detection and Constraints: The CBS Protocol detects inter-agent collisions in a CT node’s solution by checking for overlaps between all pairs of agents’ space-time volume. If a collision is detected between two agents $A,C$, a representative space-time point $(p_{0},t_{0})$ in the overlapped region is selected (in 2D this corresponds to a point $p_{0}=(x_{0},y_{0})$). To resolve this collision, we add a $(\text{agent\_id},p_{0},t_{0})$ constraint for each colliding agent that forbids the agent from overlapping spatially with $(p_{0},t_{0})$ (we note prior work [li2019largeAgentMAPF] introduced this constraint for large agents in gridworld MAPF). We highlight that the constraint is defined in the workspace dimension (e.g., 2D or 3D) and is independent of the dimensionality of each agent / solver.
A subtle but important caveat is that in continuous time motion planning, a constraint at a specific instantaneous time could be satisfied by planning a near identical path with an almost instantaneous wait, which will result in another collision at the same location at a negligibly offset time. Likewise for continuous space motion planning, a constraint at a single point could be satisfied by planning a near identical path with an almost negligible deviation. Thus, for practical purposes, the CBS Protocol should use constrain space-time volumes (e.g., forcing the agent to avoid $p_{0}$ from $(t_{0},t_{0}+\delta)$ where $\delta$ is the constraint duration) that force meaningful changes in agents’ paths. We note that there are extensive improvements in the CBS literature for designing better constraints [shaoul2024generalized_ecbs, srli2021, andreychuk2021improving], but this is not related to our main contribution and should be incorporated in the future.
Independent Tasks: CBS for motion planning has been primarily used to plan collision-free start-goal paths. From the API perspective, however, the API simply needs to find a solution that satisfies the constraints. A few prior works have used this flexibility in CBS for planning paths that visit multiple goal locations [ren2022cbs_steiner]. Thus more generally, the API is not limited to planning start-goal paths and can be used for arbitrary single-agent tasks (e.g., coverage, exploration).
@@
Theoretical Properties of the CBS Protocol: The theoretical guarantees of using the CBS Protocol depend on the single-agent APIs and the constraints used. The CBS Protocol is “complete” (i.e., guaranteed to eventually find a valid solution in a solvable instance) given two conditions: 1) Each single-agent API will find a valid solution with bounded cost that satisfies constraints if a solution exists, and 2) The constraints are mutually disjunctive. The proof of completeness follows regular CBS [sharon2015cbs].
However, as mentioned earlier, for practical performance constraints should be on space-time volumes, which are not mutually disjunctive. More broadly, in continuous time planning, designing meaningful mutually disjunctive constraints for CBS is an unsolved question [li2025cbscontinuoustimerevisit]. Thus, the CBS Protocol is not theoretically complete in the continuous space-time context, however, it still offers a principled way to resolve inter-agent conflicts for efficient motion planning.
IV Experiments